10. (8 points) Find the equation of the plane that passes through the points (1,-1, 1), (2,0,1) and (4,1,2)
Answer #1

в ( 4, 1,2). (c--) A (1,-1,2) (2,р,L) At = (x-1) + (9+1) + (-1) At = (2-1) + (+1)j +(1-1) AC = +] |-- Г AB — AB (+-1) t + (L+